OpenStudy (anonymous):

Express the required measurement in exact form. A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground?

OpenStudy (anonymous):

First you want to find the circumference of the wheel: \[C = 2 \pi r, r = 26\] Then, you only want a portion of the circumference, only 80 degrees of it, so multiply your answer by: \[\frac{80}{360}\]

OpenStudy (anonymous):

5.77

OpenStudy (anonymous):

hmm...should probably be a little bigger than that. In your calculator you want to do: \[2*\pi*26\] Then, whatever that answer is, you want to multiply it by 80, then divide by 360.